OpenStudy (anonymous):

y varies indirectly as twice x. When x = 1, y = 3. Find y when x = 4.

OpenStudy (anonymous):

x=1/2y+c (c is a constant) 1=1/(3*2) + c c=5/6 4=1/2y +5/6 1/2y=(4-5/6) 2y=1/(4-5/6) y=1/(2(4-5/6))